# Richard A Berk

### v2

@@ 11 Summary and Conclusions
There are several defining features represented in our attempt to forecast extreme heat waves. To begin, within supervised learning traditions used for forecasting, the main response variable is a binary label indicating whether an extreme heat wave has occurred or not. With so many extant definitions of heat waves, we settled on the prospect of a heat dome despite its dynamic properties. There is some evidence that extreme heat waves are more than ordinary warm summer weather (Li et al., 2024).
@@
We rely on the AIRS remote sensing data. These data do not depend on climate simulations of any kind including downscaled global climate models. AIRS data are probably noisier than simulation data and provide fewer variables with which to work; the data have promise but offer special challenges. Despite the challenges, the data in the PNW and in the Phoenix area appear to deliver instructive results.
@@
Although our work is informed by the climate science literature, we make no explicit use of models. Our analyses rest on quasi-experimental designs, a supervised learning algorithm well suited for forecasting, and a genetic algorithm for simple feature engineering. Accurate forecasting is our primary goal. Any explanatory results are a secondary benefit.
By taking the forecasting especially seriously, forecasting uncertainty had to be taken especially seriously as well. We capitalized on recent developments in nested conformal prediction sets that can provide guarantees of valid results even in finite samples as long as the calibration data are exchangeable. We make the case for exchangeability within a particular crossover design. The split sample approach has the added benefit of working with a wide range of classifiers even if they are less than ideal by themselves.
@@
We forecast extremely well for the Seattle area heat dome in late June of 2021 and for the Phoenix area heat dome in early July of 2023. The same predictors do well in both regions. Although the partial dependence plots are not identical in the two locations, a case can be made that they are qualitatively similar despite important differences in local climate, topography, research designs and data quality. And both sets of partial dependence plots also tend to be highly nonlinear. Still, more study sites surely are needed.
Finally, even if our findings are robust, there are many steps between hand-tailored research procedures and production quality forecasting tools that can be routinely used to deliver accurate heat dome warnings sufficiently in advance. We do not discuss such matters in this paper. That would be premature.

### v3

@@ 11 Summary and Conclusions
Our efforts to provide a foundation from which to forecast extreme heat waves were guided by several key decisions. To begin, within the supervised learning traditions used for forecasting, the main response variable used in training was a binary label indicating whether an extreme heat wave had occurred or not. With so many extant definitions of heat waves, we settled on a definition emphasizing a heat dome and responses to extreme heat waves that are often dichotomous: preparations to open public buildings with air conditioning as “cooling centers,” having first responders ready to check on elderly people often living alone, making preparations to close public schools in the afternoons giving ample time for parents to arrange for child care, requiring for water breaks at outdoor work sites, re-allocating staff in emergency rooms to handle increased patient volume, and stocking emergency rooms and urgent care centers with adequate medical supplies (e.g., IV fluids, electrolyte solutions, ice packs, heat-related medication), and much more.
@@
We relied on the AIRS remote sensing data. These data do not depend on climate simulations of any kind including downscaled global climate models. This seems especially important because of our focus on rare, extreme events. AIRS data are probably noisier than simulation data and provide fewer variables with which to work; the data have promise but offer special challenges. Despite the challenges, the data in the PNW and in the Phoenix area appear to deliver instructive results.
@@
Although our work has been informed by the climate science literature, we make no explicit use of models. Our analyses rest on quasi-experimental designs, a supervised learning algorithm well suited for forecasting, and a genetic algorithm for simple feature engineering. Accurate forecasting is our primary goal. Any explanatory results are a secondary benefit.
@@
The classification in the Pacific Northwest and in the greater Phoenix area proved to be extremely accurate. Although the variable importance plots and partial dependence plots were not identical in the two locations, a case can be made that they are qualitatively similar despite important differences in local climate, topography, research designs and data quality. Still, many more study sites surely are needed, ideally with many more grid cells per site.
By taking genuine forecasting seriously, forecasting uncertainty had to be taken seriously as well. We capitalized on recent developments in nested conformal prediction sets that can provide guarantees of valid results even in finite samples as long as the calibration data are exchangeable. However, the manner in which the data were curated violated exchangeability. Future work can put true forecasting in play by not conditioning on the particular days in which extreme heat waves occured. An illustration, whose details would need to be worked out, could use the AIRS data and a trained classifier perhaps of the sort trained on the PNW data. One could consider each day in turn for the PNW summer months of June, July and August of 2025 (assuming AIRS data availability). For each day, the observed predictor values would be input to the trained classifier. On days when the predicted probability for an extreme heat wave 2 weeks later was sufficiently high, an extreme heat wave would be forecasted. The definition of “high” would depend on the predetermined tradeoff of false positives to false negatives. Conformal prediction sets might be valid as well. There would be much to think through, but no conditioning on a known heat wave day.
We acknowledged the issue of endogenous sampling, but argued that corrections for response variable imbalance are only part of the problem. The cost ratio of false negative and false positive classification errors is implicated as well. Sample imbalance and misclassification cost ratios are mathematically linked and could imply the need for an asymmetric loss function. We briefly considered some other options but currently lack the requisite information from a real policy setting having engaged stakeholders. To start the discussion, our results are based on within-subject designs that achieve adequate sample balance and a cost ratio of about 1 to 1.
Even if our work proves useful, it is at best a start. Among the more important issues going forward is our choice of a two week lag between forecasting precursors and heat wave outcomes. We relied on recent research, but a two week lag is hardly definitive. Our extreme heat wave definition also needs to be revisited. Empirically, it seems to work well, but others might work well too.There are many options already in print. Another significant matter is formulating appropriate populations that are instructive and yet provide better balanced data. We suspect that even if promising populations are proposed, endogenous sampling complications will be ongoing. Extreme heat waves may be increasing, but they are still relatively rare, and “compared to what” questions will remain. Also remaining are research designs and analysis that undertake valid forecasting. Finally, there are many steps between hand-tailored research procedures and production quality forecasting tools that can be routinely used to deliver accurate heat dome warnings sufficiently in advance. We did not address such matters here. That would be at least premature.
Acknowledgements Eric Tchetgen Tchetgen offered very insightful concerns about endogenous sampling.